9) Chemical Formula: C5H1002 IR: strong peak at 1750cm-1 6H PPM10) Chemical Formula: C12H16 (hard to see but, all peaks that integrate to 1H are doublet of doublets) IR: peak at 1650cm-1 9

Please explain how to read and draw the Chemical structure for the following questions depending on the information provided: Chemical Formula, IR and NMR. I really want to understand how to solve these.
Answer #1

C5 Ho0 CDBE 5 -e z 6-5=1) S20C513H) S5.0 Cseptt,i carbon stdtch ester C 1HIC CPBE 2 -16 = u+1) =(3 - F3 CS9H S1.Cd S.LS Cddau